OpenStudy (anonymous):

Please Help Will Medal! solve in 2 +in x=5 round to the nearest thousandth is necessary

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[\ln(2)+\ln(x)=5\\ \ln(2x)=5\] is the first step